40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 30 13:02:51 2013, Masayuki,Manasa, Update, IOO, PMC and MC alignment 
    Reply  Thu Oct 3 01:24:31 2013, Jenne, Rana, Update, IOO, PMC  PMCR_1064822387.bmppmc80.png
       Reply  Sun Oct 6 22:52:09 2013, Jenne, Rana, Update, IOO, input beam to PMC aligned again pmcr.png
          Reply  Mon Oct 14 17:51:40 2013, Jenne, Update, IOO, input beam to PMC drifted again 
             Reply  Mon Oct 21 13:15:52 2013, Koji, Update, IOO, PMC aligned 
             Reply  Fri Oct 25 10:45:16 2013, Steve, Update, PSL, laser drift monitor set up laserDriftMon.jpgPMCT_120d.pngPMCT_1000d.png
                Reply  Fri Oct 25 19:56:58 2013, rana, Update, PSL, laser drift monitor set up PMC-IRSISSS.pngPA250052.JPGPA280044.JPG
                Reply  Fri Jan 10 15:33:02 2014, Steve, Update, PSL, laser drift monitor set up idea 2daysDrift.png
                   Reply  Mon Jan 13 11:08:48 2014, Steve, Update, PSL, 3 good days of IOO pointing 3gdPSLpointing.png
                   Reply  Tue Jan 14 10:12:12 2014, Steve, Update, PSL, laser drift monitor set up idea bad4thday.png
Message ID: 9549     Entry time: Mon Jan 13 11:08:48 2014     In reply to: 9547
Author: Steve 
Type: Update 
Category: PSL 
Subject: 3 good days of IOO pointing 

 Three good days of IOO pointing: Friday, Sat and Sun    What was changed?  May be the clamping on Friday?

IOO vertical changes recovering as tempeture. IP is clipping at plastic enclosure of ETMY

 

NOTE: ANTS at the PSL optical table.  I will mop with chemicals tomorrow if we see more.

 

Attachment 1: 3gdPSLpointing.png  46 kB  | Show | Show all
ELOG V3.1.3-